### v2

# ALMA chemical survey of disk-outflow sources in Taurus (ALMA-DOT)

Abstract
Context. The ALMA image of the ringed disk around HL Tau stands out as the iconic signature of planet formation, but the origin of the observed substructures is still debated. The HL Tau system also drives a powerful bipolar wind, detected in atomic and molecular lines, that may have important feedback on the process.
Aims. The outermost component of the wind traced by CO emission is analyzed in detail to determine its relationship with the disk and its substructures.
Methods. A spectro-imaging investigation is conducted using ALMA observations of the 12CO (2-1) line at 1.3 mm, with 0.2 km sand spectral and angular resolution, in the framework of the ALMA-DOT project. Relevant wind parameters are derived, allowing a tomographic reconstruction of the morphology and kinematics of the red-shifted lobe of the outflow to be compared with theoretical models.
Results. The data channel maps and position-velocity diagrams show a rich substructure of concatenated bubble- and arc-shaped features, whose size and distance from the source continuously increase with velocity. The superposition of such features generates the apparent conical shape. The spatio-kinematical properties suggest that the flow presents distinct nested shells with increasing velocity and steeper velocity increase with distance from the star going toward the axis, and rotating in the same sense of the disk. The wind parameters are compared with the predictions of magnetohydrodynamic (MHD) disk winds. Under this hypothesis, the launch radii of the three outermost shells are found to cluster around the positions of three rings in the dust distribution of the disk located at 58, 72 and 86 au from the star. We derive a magnetic lever arm , higher than that commonly adopted in models of MHD winds from the outer disk. Interpretations are discussed.
Conclusions. The properties of the CO outflow from HL Tau appear to be compatible with magnetized disk winds with launch radii in the region at 50 to 90 au from the source. As such, the wind may be capable of removing angular momentum also from the outer disk. The arrangement of the wind in nested shells with brighter emission rooted at the location of ring substructures could support the results of recent non-ideal MHD simulations according to which magnetic instabilities can spontaneously generate both the ring-gap system and a connected inhomogeneous layered wind, alternatively to the action of yet undetected protoplanets. Further observational analyses and comparisons with other classes of models will help establish the role of magnetic effects in the process of planet formation.
Key Words.:
Protoplanetary disks – Jets and outflows – ISM: molecules – Stars: individual: HL Tau1 Introduction
One of the most spectacular discoveries of the last decade has been the detection in the continuum emission at millimeter wavelengths of concentric gaps and rings in disks around young stars. The first example found is the disk around HL Tau. Observed with the Atacama Large Millimeter Array (ALMA)111https://www.eso.org/sci/facilities/alma.html, it showed in exquisite detail a sequence of concentric rings and gaps (ALMA Partnership et al., 2015), now commonly seen in many other disks (e.g., Andrews et al. 2018).
The first interpretation offered was that these gaps are carved by young massive planets embedded in the disk (e.g., Lodato et al. 2019). However, this scenario raises open issues, such as the amount of mass necessary to produce massive planets at such large distances from the star (up to 100 au in HL Tau) and in such short timescales (less than 1 Myr). In addition, the predicted embedded planets inside the gaps remain elusive so far, except for PDS 70, where the ”gap” actually is a large inner cavity (Keppler et al., 2018).
Several other disk processes are able to produce ring/gap structures. Perhaps the most generic are magnetic instabilities in the low-ionization ”dead zone” (e.g., Suriano et al., 2018; Riols & Lesur, 2019). This process could be widespread in young disks that harbor a residual primordial magnetic field. In this case, an inhomogeneous magnetohydrodynamic (MHD) disk wind (DW) is created, and the disk self-organizes in a system of rings and gaps, with denser wind layers launched from the rings and less dense ones from the gaps (e.g. Suriano et al. 2019; Riols et al. 2020). An asymmetry also develops between the red and blue lobes (Béthune et al., 2017), as observed in almost all cases (e.g. Podio et al. 2011, 2021). Finding firm evidence for such layered MHD disk winds is thus an essential step in a clarification of the origin of disk substructures and of the role of magnetic fields in planet formation.
Signatures of layered outflows have indeed been found with ALMA in a growing number of young systems, in the form of rotating molecular winds flowing in a cavity of apparent conical shape (e.g., TMC1A, (Bjerkeli et al., 2016); HH 46/47, (Zhang et al., 2016, 2019); HH212 (Lee et al., 2017); DO Tau, (Fernández-López et al., 2020); HH 30, (Louvet et al., 2018; López-Vázquez et al., 2024; Ai et al., 2024); DG TauB (Zapata et al., 2015; de Valon et al., 2020, 2022); HH270mms1-A, (Omura et al., 2024); L1448-mm, (Nazari et al., 2024)). The CO outflow is associated in some cases with a coaxial molecular wind that emits in lines, with a fast collimated atomic jet flowing along the axis of the system (e.g., HH 46/47, (Nisini et al., 2024), DG Tau B (Delabrosse et al., 2024)). The origin of the layered structure in these CO winds is debated. In addition to the magnetic instabilities mentioned above, other studies invoke the formation of swept-up wind shells (Zhang et al., 2019), peculiarities of the MHD disk-wind accretion/ejection engine (Louvet et al., 2018; de Valon et al., 2020), flow wiggling (de Valon et al., 2022), the action of the wings of subsequent bow shocks of time-variable jets (Tabone et al., 2018; Rabenanahary et al., 2022), or the evolution of a unified wide-angle magnetized flow with a jet-bearing X-wind (Shang et al., 2020, 2023; Ai et al., 2024).
Among the observable sources, HL Tau stands out as the first and only object where a possible coincidence between the wind structure and the distribution of disk ring/gaps can be studied. In fact, in the other cases the flow is associated with a disk where no gaps and rings are detected. In some cases, this can be due to the geometry of the observation (for example, the edge-on disks of HH 30). In general, however, the accretion/ejection activity associated with young stellar objects is expected to decrease with the source evolutionary stage; hence powerful and bright molecular winds and jets are associated with Class 0 protostars (e.g., Lee et al. 2018; Tabone et al. 2017; Podio et al. 2021) while only faint molecular emission is associated with evolved Class II stars. On the other hand, substructures such as rings and gaps are commonly detected in pre-main sequence disks (e.g., SHARP LP, Andrews et al. 2018), while they are rarely detected in younger disks (see, e.g., Segura-Cox et al. 2020; Sheehan & Eisner 2017, 2018; de Valon et al. 2020). In this context, HL Tau is intermediate between an evolved accreting protostar (Class I) and a classical TTauri star (Class II), and is associated with both a disk with clearly detected rings/gaps and a bright extended and structured outflow. HL Tau is therefore an ideal target for investigating the possible relationship between the disk substructures and the wind properties and for testing the proposed scenarios for their formation.
The central object, a young star of mass 2.1 M⊙ (Yen et al., 2019a), is the source of a composite outflow seen at optical, infra-red, and mm wavelengths. The collimated atomic jet that shines in shock-excited lines such as [S II]6716,6731, H, and [O I]6300, stretches out of the reflection nebula surrounding the star for about 1.5 along the direction PA 50°, with an inclination angle with respect to the line of sight of (Mundt et al., 1990; Anglada et al., 2007; Pyo et al., 2006; Movsessian et al., 2012). The jet is compatible with being perpendicular to the disk, which has an inclination angle with respect to the line of sight = 46.2 (defined as for face-on), as determined in ALMA Partnership et al. (2015) from the data at 02 resolution. The same study reports that the major axis of the disk lies along PAdisk = +138.2 measured from north to east.
The system was imaged at 02 resolution with adaptive optics in the near-infrared, showing a collimated jet in [FeII] 1.64 m and a surrounding wider angle warm wind emitting in H2 2.12 m , that flows in a conical cavity detected in scattered light (Takami et al., 2007; Beck et al., 2008). A coaxial molecular outflow has been observed in the CO(3) and CO(1) lines (Lumbreras & Zapata, 2014; ALMA Partnership et al., 2015; Klaassen et al., 2016). The data show a prominent conical redshifted flow to the south-west of HL Tau, with the axis almost aligned with the atomic jet, and with a full opening angle starting at 90°and then narrowing at 60°further from the source (Klaassen et al., 2016). The inclination with respect to the line of sight is assumed to be along the perpendicular to the disk (i.e., =). The north-eastern lobe of the molecular outflow is much fainter and appears as a wide cavity spread within the ambient.
In this paper, we investigate the structure of the CO outflow in more detail analyzing a dataset obtained within the ALMA-DOT program (ALMA chemical survey of Disk-Outflow sources in Taurus), an observational campaign aimed at characterizing the gas in disks and outflows of embedded sources in the Taurus star-forming region. The high angular/spectral resolution of ALMA allows us to disentangle and characterize the emission from the various elements of the systems (disk, outflows, streamers, envelope). Previous papers in the ALMA-DOT series focus on the chemical properties of the disks (Podio et al., 2019, 2020a, 2020b; Codella et al., 2020) and of the associated streamers and outflows (Garufi et al., 2020, 2022), while Garufi et al. (2021) describes the collective properties of the targets examined as a whole. The present work focuses on the observations of the molecular outflow around HL Tau, as imaged in the line of 12CO (2) at 1.3 mm with a resolution of 03 and 0.2 km s, and on the implications of these findings for the structure of the disk.
The paper is organized as follows. Section 2 details the observations and the reduction of the data. Section 3 describes the results obtained, while the derived hypothesis for the structure of the wind is described in Section 4. In Section 5 we discuss the comparison of the results with theoretical models in an MHD disk wind scenario. The conclusions of this study and the perspectives for future analyses are outlined in Section 6.
2 Observations and data reduction
Cycle 6 observations of the ALMA-DOT campaign (Podio et al., 2019) were taken with ALMA-Band 6 on October 28, 2018 in an extended configuration with baselines ranging from 15 m to 1.4 km (project 2018.1.01037.S, PI: L. Podio). The field of view is a circular region around HL Tau of radius 185, corresponding to about 2.7103 au at the distance from the source, 147.3 pc (Galli et al., 2018). The integration time sums up to 110 minutes (Garufi et al., 2021). The bandpass and phase calibrators were J0423-0120 and J0510+1800, respectively. The correlator setup consisted of high-resolution (0.141 MHz) spectral windows (SPWs) covering 12CO (2), CN (2), o-H2CO (3 CS (5), and CH3OH () (A). The molecular parameters were taken from the Cologne Database of Molecular Spectroscopy (CDMS) (Müller et al., 2001). In this paper, we focus on the 12CO (2) emission (rest frequency 230.538 GHz), while the description in other molecular lines can be found in Garufi et al. (2020).
Data were reduced using the standard procedure using the Common Astronomy Software Applications package (CASA, McMullin et al. 2007) version 4.7.2. Self-calibration was performed on the continuum emission and applied on the line-free continuum, and the continuum subtracted line emission. The procedure improved the continuum S/N by a factor of 3.3. The final continuum maps and spectral cubes were produced with tclean by applying a manually selected mask to the signal and iterating until the residuals revealed no significant source emission. We adopted Briggs weighting with robust=0.0 for the bright CO line, to maximize angular resolution, and set a channel width of 0.2 km s. The clean beam of the self-calibrated maps in CO is 0.26″0.31″ with a PA of -3.12°. The r.m.s. noise per channel is mJy beam-1. The beam in the continuum images is 0.23″0.26″ with PA= -6.77°. The r.m.s. for the continuum image is 0.12 mJy beam-1.
3 Results
3.1 Continuum emission
The continuum at 1.3 mm has an integrated flux of 844.9 1.9 mJy, with peak intensity of 95.48 0.19 mJy beam-1 located at R.A. 04h 31m 38s.43, and Dec. 18∘ 13′ 5717. At our angular resolution the intensity distribution is smooth, and the sequence of rings and gaps is not visible. The FWHM along the major and minor axes, determined with a 2D Gaussian fit deconvolved from the beam, is and mas, respectively. From these values, we estimate an inclination of the disk (with 0° corresponding to face-on). The fit gives a disk position angle, PAdisk = +138.25. Both values agree with those reported in ALMA Partnership et al. (2015).
3.2 CO (2) emission
The ALMA datacube in the CO (2 line consists of 290 velocity channels with VLSR ranging from -26 to +32 km s, in steps of 0.2 km s. The emission of the outflow is detected from -10 km sto +31.8 km s. The systemic velocity is Vsys=+7.1 km s(Garufi et al., 2021).
Figure 1 presents the intensity integrated over the velocity (moment-0 map) in the full field of view. The left panel shows the emission integrated over the blue-shifted velocity range V, while the integration in the right panel is for VLSR velocities between +8 and +30 km s, illustrating the red-shifted emission. The faint blue-shifted emission is distributed along a wide-angle cavity of semi-spherical shape pointing towards the north-east (NE), with a distinct bright feature on the northern side. The emission in the red-shifted south-west (SW) lobe presents two main components, an inner flow of apparent conical shape around an axis oriented at PA = 228°(coinciding with the minor axis of the disk) with full opening angle starting off at 90°and then narrowing at 60°further from the source (Klaassen et al., 2016) and an outer faint conical wide wind, surrounding in projection both the disk and the base of the inner outflow, with vertex at 15 NE of the source and opening angle of 120° . To guide the eye, the approximate spatial limits of these components are indicated by solid white curves in the figure.222 The opening angle was determined with a linear fit to the wind edges, identified by the position of the zeroes of the second-order derivative of the emission along cuts parallel to axis, in the first 2″of propagation. A molecular emission ridge is evident at 14” from the source. This is part of a wider elliptical feature, as shown in ALMA Partnership et al. (2015), Fig. 1, and reported here with a dashed white elliptical curve.
The intensity-weighted mean velocity (moment-1 map) of the emission is presented in Fig. 2. The NE sector hosts blue-shifted material that is not collimated, while the red-shifted SW lobe is structured. The transition in velocity between the two wind components of the red lobe identified in the moment-0 map shows that the outer wide wind is slower than 10 km s, while the inner flow has increasing velocities from the border toward the axis, up to about 30 km s.
The asymmetry between the blue- and red-shifted lobes may be an effect of the inclination angle with respect to the plane of the sky. For example, Klaassen et al. (2016) illustrates how the large aperture of the initial portion of the HL Tau flow causes the far side of the approaching lobe to be imaged at red-shifted velocities close to the systemic. However, while this effect may explain the loss of the upper part of the large arc at blue-shifted velocities, it would require the approaching lobe to have an opening angle larger than the receding one to explain the absence of a well-developed conical structure. A trace of faint conical arms of aperture angle similar to that in the red lobe is barely seen at the base of the blue-shifted lobe, but it is not seen further out, and it is covered by the signal of the large semicircular arc. This structure of the blue lobe may be due to HL Tau being located on the wall of a large-scale molecular bubble inflated by nearby XZ Tau, at 23″east of the source (Welch et al., 2000; Yen et al., 2019b). The expanding bubble may have compressed the outer envelope of HL Tau on its NE side, smashing the CO cavity, while the SW lobe remained unaffected. Signatures of such a large-scale bubble are visible at low blue-shifted velocities in Fig. 2.
In the following we concentrate the analysis on the red-shifted lobe that presents a well-defined and relatively bright structure over a wide range of velocities. This allows a detailed inspection of the spatio-kinematic properties, for a meaningful comparison with proposed acceleration mechanisms. We defer the study of the fainter blue-shifted lobe to a forthcoming investigation, adding that the squared 10″at the center of the region were discussed in Garufi et al. (2022), where the kinematics of the rotating disk is analyzed from the same dataset, as well as possible accretion streamers.
3.3 Velocity channel maps
A comprehensive atlas of the channel maps of the CO (2-1) emission is given in Appendix A. The analysis of the individual channels reveals a faint but very rich flow substructure, not visible in the moment-0 and moment-1 maps. The emission in the blue lobe is barely detectable, but the signal is strong in the red-shifted lobe, allowing for detailed inspection. A selection of channel maps illustrating the configuration of the inner flow in the red-shifted SW lobe is presented in Fig. 3. The maps show an apparent conical structure, with aperture decreasing with velocity, from about 80 at 10 km sto 55 at 26 km s(see Appendix A).
Inscribed in the conical apparent shape, one finds a series of ‘bubbles’ and ‘arcs’ of different sizes, positions and brightness. The arcs appear to be part of concatenated closed curves that can be visually fitted by ellipses, as illustrated in Fig. 4. About ten such structures can be identified, grouped in four families that we label as A, B, C, D. Each feature continuously changes size and position with channel maps and acquires progressively larger extension and separation from the star as VLSR velocity increases, but maintains its shape unchanged. Arcs ’A’ are observed at low velocity (10-16 km s, see Fig. 4a-b) and at small separation. Despite the blending of the emission on the source side and the fragmented appearance at larger distances, these structures clearly form closed bubble-like curves expanding with increasing velocities. Arcs B1 and B2 have a large extension and only show the convex branch on the side of the source. They can be followed over a wide range of velocities, from 11 to 19 km s, in which they continuously change position from 3 to 10 from the source, as illustrated in Fig. 17. Arcs C are visible only at high velocity (20-29 km s, see Fig. 4c-d. They are relatively faint, but their positional shift with velocity can be clearly recognized. Finally, the bright arc D is visible at a small distance from the source (about 09) and at intermediate velocities (20 km s, see Fig. 4c. The orientation of the features deviates slightly from that of the axis of the system. This is illustrated, for example, in panel b of Fig. 4.
In general, inspection of channel maps makes clear that the flow cannot be defined to have a simple and continuous conical shape. In contrast, the apparent conical border is revealed to arise from the superposition of the discrete limb-brightened arc-shaped substructures.
3.4 Parallel Position-Velocity diagram
Figure 5 illustrates the Position-Velocity diagram obtained with a virtual slit 03 wide (comparable to the beam size) positioned along the axis (PA=228∘). The diagram (hereafter referred to as PV∥) collects emission from both circumstellar material and extended wind. The emission within offset 2 arcseconds from the source is due partly to the disk and partly from the base of the wind, whose large opening produces contributions in both red-shifted and blue-shifted velocities on each side (e.g., Klaassen et al. 2016). The emission at larger offsets and velocities comes only from the extended wind.
The striking characteristic of the diagram is the fact that the emission in the wind region is not continuous, but it is distributed in a fan of separated traces on both sides of the source (superposed to a low intensity more homogeneous emission in the region and VLSR km s). In the blue lobe, at least three linear traces are seen to propagate towards larger blue-shifted VLSR velocity. On the SW red-shifted side, more than ten traces can be recognized. Within the first 4″a fan of almost linear traces opens from the source reaching high velocities, up to at least VLSR 28 km s. For each of them, the velocity appears to increase linearly with offset, in a way reminiscent of a ’Hubble law’ trend. This group of traces displays the steepest slope, that is, the highest apparent acceleration, among the features in the diagram. A second fainter fan of almost linear segments is observed between 4 and 11″, with lower VLSR velocities up to 15 km s. Interposed between the two groups, one finds two extended traces of curved shape, which reach velocities as high as 20 km s.
The substructures detected in the diagram suggest a common origin with the arc- and bubble-shaped features identified in the channel maps. To illustrate this hypothesis, for each of the ellipses described in Section 3.3, and for each velocity channel, we report on the red-shifted portion of the PV∥ diagram the offset of the intersections of the fitted ellipses with the flow axis (see Fig. 4). The subscripts ’’ and ’’ indicate the intersections of the side of the ellipse closer and farther from the source, respectively. The broken lines obtained by connecting the points relative to the same side of each ellipse are strikingly coincident with sections of the emission branches in the diagram.
The first group of traces reaching high velocities corresponds to arcs D (orange), C (red) and to the intersections of type ’’ of features A (green). The second group corresponds to the SW sides of the A ellipses, marked as ’’ (green). Given the inclination angle of the wind in space (=226∘.5), the A0 traces appear to correspond to the rear side of the flow and the A1 traces to the front side. Note that the A0 and A1 traces appear in the same order from the source, in a way that is consistent with the fact that in the channel maps the ellipses of the A family are concatenated rather than nested. The long curved traces correspond to the features of the B family (in blue). These filaments well reproduce the consistent shift in position with velocity illustrated in Fig. 17. The emission peak at offset 14 at a constant low velocity named ’E’ corresponds to the ridge of emission south-west to the source and almost perpendicular to the flow direction seen in Fig. 1.
3.5 Transverse Position-Velocity diagrams
Appendix B presents an atlas of the transverse Position-Velocity diagrams (hereafter PV⟂) obtained for the red lobe with a 03 wide virtual slit centered on the axis and oriented at PA=318° (i.e. perpendicular to the axis and with positive offsets toward NW). The diagrams are formed at the separation of 35 to 10″ from the source, in steps 05. Illustrative examples are given in Fig. 6, for selected separations from the star.
In all panels, the emission presents a bell shape around the axis, narrower toward high velocities. Progressing with separation from the source, the bell becomes wider, but the lateral ’walls’ maintain coherence. Outside of the bell, only low-velocity material (7.5 - 9.5 km s) is present.
Internal structures in the form of arcs and branches of ellipse-shaped curves are found in all the panels, with a monotonic increase in average velocity and size as the separation from the source increases. This behavior is reminiscent of the variation with velocity observed for the substructures in the channel maps and in the PV∥ diagram. Despite the additional difficulty due to the signal confusion created by the overlap of the emission traces, Figure 6 proposes an identification of visible curves with features belonging to the A and B family, with the same labels as in Fig. 5.
An evident characteristic of the structures identified in the transverse diagrams is the skew with respect to the horizontal. This property is related to the rotation of the flow. The two opposite borders of the wind, in fact, present a velocity shift due to the projection of the rotation (i.e. toroidal) velocity along the line of sight. An illustration of this effect is offered in de Valon et al. (2022), Fig. 2. The observed features display higher red-shifted velocities on the NW side (seen more clearly for the curves of the B family) consistent with a clockwise rotation sense looking from the tip of the red lobe toward the source. This sense of rotation matches that of the disk (Garufi et al., 2021).
In summary, the ensemble of maps presented in this Section indicates that the orthogonal projections of the Position-Position-Velocity (PPV) datacube all show well-defined substructures, consisting of concatenated curves in the velocity channel maps and in the transverse PV⟂ diagrams, and of a fan of traces opening from the source region in the parallel PV∥. The features can be connected from one projection to the other, and this leads to inferring the existence of coherent structures in the datacube.
4 Interpretation of the substructures in terms of nested gas shells.
In this section, we analyze what type of wind could determine the emission properties in the PPV datacube. The description above suggests a wind configuration in physical space consisting of a series of nested rotating shells, which gradually open up with distance from the source, with a progressively higher propagation speed and a steeper speed gradient with distance from the source for the shells closest to the axis. An illustration of this hypothesis is provided in the sketch of Fig. 7, based on an axisymmetric flow composed of three such layers, which is assumed not to rotate for simplicity.
In the real case, rotation and inclination introduce deformations of the traces in the projected maps. The skew of the features due to flow rotation has been described in the previous Section. On the other hand, inclination with respect to the line of sight produces a deformation of the closed curves with respect to the pure ellipse shape, as described, e.g., in De Valon (2021). We adopt for the inclination angle of the red-shifted flow with respect to the line of sight a value 46.5° . This compares well with the inclination angle of the red-shifted axial atomic jet, , determined by combining the tangential velocity measured for the most coherent [SII] emission knots on the SW side of the source in Anglada et al. (2007) with the average radial velocity measured for the same jet lobe by Mundt et al. (1990). Due to this inclination, the low-velocity part of the ellipses is expected to be flatter than the high-velocity part. However, this deformation is not easily recognizable in our maps, since the structures centered on moderate velocities are weak and/or do not have a visible upper part, while those centered on low velocities have their lower part possibly hidden by the absorption of CO gas with velocities close to Vsys.
In any case, the overall appearance of the data supports an interpretation of the observed substructures in terms of distinct rotating shells nested in physical space. As mentioned in the Introduction, the same hypothesis has been put forward for other molecular winds that exhibit similar spectro-imaging properties in high-resolution data.
However, it is important to note that HL Tau is the only object among the known cases with ring / gap substructures detected in the disk. In this case, therefore, it is possible to study the relationship between the substructures observed in the wind and those visible in the disk.
4.1 Tomographic reconstruction of the wind shells.
[arcsec] | [arcsec] | [arcsec] | [km s] | [km s] | [103 au] | [102 au] | [km s] | [km s] | [deg] | [km s] | [102 au km s] | [au] | ||
B1a | 4.5 | -3.50 | 3.54 | 11.95 | 14.68 | 0.9 | 5.2 | 9.0 | 1.88 | 21.1 | 9.7 | 9.8 | 57 | 3.0 |
5.0 | -3.88 | 3.84 | 12.40 | 15.10 | 1.0 | 5.7 | 9.7 | 1.86 | 21.1 | 10.4 | 10.6 | 58 | 3.2 | |
5.5 | -4.06 | 4.06 | 12.82 | 15.47 | 1.1 | 6.0 | 10.2 | 1.83 | 21.1 | 11.0 | 10.9 | 56 | 3.4 | |
6.0 | -4.35 | 4.27 | 13.25 | 15.95 | 1.2 | 6.3 | 10.9 | 1.86 | 21.1 | 11.7 | 11.8 | 57 | 3.6 | |
6.5 | -4.59 | 4.57 | 13.41 | 16.08 | 1.3 | 6.7 | 11.1 | 1.84 | 21.1 | 11.9 | 12.4 | 58 | 3.8 | |
7.0 | -4.84 | 4.95 | 13.46 | 16.11 | 1.4 | 7.2 | 11.2 | 1.83 | 21.1 | 12.0 | 13.2 | 61 | 3.9 | |
B1b | 4.5 | -3.89 | 3.67 | 10.80 | 13.40 | 0.9 | 5.6 | 7.3 | 1.79 | 23.2 | 7.9 | 10.0 | 71 | 2.8 |
5.0 | -4.24 | 4.02 | 11.34 | 13.89 | 1.0 | 6.1 | 8.0 | 1.76 | 23.2 | 8.7 | 10.7 | 69 | 3.0 | |
5.5 | -4.50 | 4.41 | 11.67 | 14.18 | 1.1 | 6.6 | 8.5 | 1.73 | 23.2 | 9.2 | 11.4 | 70 | 3.2 | |
6.0 | -5.02 | 4.53 | 11.90 | 14.50 | 1.2 | 7.0 | 8.9 | 1.79 | 23.2 | 9.6 | 12.6 | 74 | 3.4 | |
6.5 | -5.15 | 4.76 | 12.09 | 14.65 | 1.3 | 7.3 | 9.1 | 1.76 | 23.2 | 9.9 | 12.9 | 73 | 3.5 | |
7.0 | -5.43 | 5.15 | 12.32 | 14.90 | 1.4 | 7.8 | 9.5 | 1.78 | 23.2 | 10.3 | 13.9 | 75 | 3.7 | |
B2 | 6.0 | -5.34 | 4.70 | 11.10 | 13.60 | 1.2 | 7.4 | 7.6 | 1.72 | 17.5 | 8.0 | 12.7 | 90 | 3.1 |
6.5 | -5.52 | 5.03 | 11.50 | 13.90 | 1.3 | 7.8 | 8.1 | 1.65 | 17.5 | 8.5 | 12.9 | 85 | 3.2 | |
7.0 | -5.70 | 5.52 | 11.60 | 14.10 | 1.4 | 8.3 | 8.4 | 1.72 | 17.5 | 8.8 | 14.2 | 89 | 3.5 | |
7.5 | -5.89 | 5.65 | 12.08 | 14.43 | 1.5 | 8.5 | 8.9 | 1.62 | 17.5 | 9.4 | 13.8 | 83 | 3.5 | |
8.0 | -6.18 | 5.73 | 12.29 | 14.57 | 1.6 | 8.8 | 9.2 | 1.57 | 17.5 | 9.6 | 13.8 | 80 | 3.6 | |
8.5 | -6.47 | 5.84 | 12.33 | 14.67 | 1.7 | 9.1 | 9.3 | 1.61 | 17.5 | 9.8 | 14.6 | 84 | 3.7 | |
9.0 | -6.56 | 6.09 | 12.41 | 14.70 | 1.8 | 9.3 | 9.4 | 1.58 | 17.5 | 9.8 | 14.7 | 84 | 3.7 | |
9.5 | -7.10 | 6.39 | 12.40 | 14.80 | 1.9 | 9.9 | 9.4 | 1.65 | 17.5 | 9.9 | 16.4 | 92 | 4.0 | |
10 | -7.16 | 6.53 | 12.45 | 14.87 | 2.0 | 10.1 | 9.5 | 1.67 | 17.5 | 10.0 | 16.8 | 93 | 4.1 |
The series of PV⟂ diagrams illustrated in Appendix B allows, in principle, a ’tomographic reconstruction’ of the wind shells that provides information on their morphology and kinematics in the physical space. This can be done by measuring the width, inclination, and position of each detected feature in adjacent diagrams.
In practice, for each PV⟂ diagram one has to identify the SE and NW apexes of the identified features and measure the corresponding position and VLSR velocity. These coordinates are sufficient to determine the local shell radius , the toroidal velocity and the axial velocity (referring to a co-moving cylindrical coordinate system centered on the star, with directed toward PA=228°with an inclination angle =46.5° ). However, the task is obstructed by the confusion generated by the faintness of the signal and the overlap of structures at low VLSR. The main difficulty lies in identifying a coherent feature that can be followed in its evolution from a diagram to the next. However, we illustrate the measurement by focusing on the features of family B, which are more clearly identified and have a regular shape. Note that a close inspection of the datacube projections reveals that the B1 feature is actually composed of two intersecting curves that we label B1a and B1b. The analysis of the other fainter arcs, requiring a dedicated procedure, will be presented in the next paper of the series (Nony et al., in prep).
To ensure the identification of the apexes in a way that is as reliable as possible, a cross-check was first performed with the positions of the emission peaks in common with the other projections of the datacube. Secondly, a mathematical fit of the sequence of peaks along the identified traces was performed, assuming an ellipse as the geometrical curve closest to the shape of the features. The SE and NW extremes of the fitted ellipses, that is, the points on the curve furthest from the axis, are considered the vertices of each feature at the shell’s edges, and their coordinates are recorded. The procedure is illustrated in Fig. 8, and details on sampling criteria and fitting routine, as well as a discussion of inherent limitations, are available in Appendix B.
Defining as () and () the coordinates of the SE and NW apexes of a given feature in the PV⟂ diagram formed at separation from the source, then:
where +7.1 km sand =46.5∘.444 In writing these formulas we neglect flow wiggling and precession. A check on the validity of this assumption is reported in Appendix C. In addition, the tangent angle to each shell is found with a linear fit to the variation of with . If the wind flows along the shell, the poloidal velocity is tangential to the surface, and . To validate this assumption, one should measure independently . Unfortunately, this measurement is complicated by the inclination of the flow with respect to the l.o.s., which distributes the information at displaced offsets (see Fig. 21), in regions where the emission traces are faint and/or confused. This makes the test conditions prohibitive in almost all cases. However, we could measure and confirm the validity of the hypothesis at least for a few positions, as described in Appendix D. In the following, it is assumed that the flow streams along the shell surfaces.
A similar but not identical tomographic analysis has been performed for the study of the flow from the young star DG Tau B in de Valon et al. (2022). The comparison of the two approaches is discussed in Appendix E.
The parameters derived from the fit for the offsets in which the B substructures are well identified are given in Footnote 3. The same quantities are illustrated in Fig. 10 as a function of both the offset and the deprojected distance from the source along the flow axis. For a visualization of the position of the three considered shells in the flow portion examined, see Fig. 9, which reports the spatial offsets of the SE and NW apexes of the ellipse fit ( and in Footnote 3) superposed to a moment 0 intensity map integrated over the velocity range 10.6 - 16.2 km s, (characteristic of the studied traces), rotated to have the axis along the vertical.
The three shells turn out to have a similar trend for the radius, that increases from about 5102 to 103 au for distances from 103 to 2103 au from the star, and appear indeed to be nested, with B1a the innermost one and B2 the external layer. The smaller opening angle of B2 with respect to the one determined for B1a and B1b at offsets closer to the source, may indicate recollimation at large distances. The axial velocity , between 7 and 11 km sat the distances sampled, is higher for the inner shells and increases monotonically. In shell B2 the increase is less pronounced after 1.6103 au, and the same occurs for B1a and B1b, but closer to the source, after 1.3103 au. The poloidal velocity , proportional to , shows a similar trend. The toroidal velocity is constant on average along a single shell, but is progressively higher toward the axis, with an average of 1.85, 1.77 and 1.64 km s, for B1a, B1b, and B2, respectively. The specific angular momentum varies between 950 and 1700 au km s. Its value increases monotonically and appears to be similar in the three shells. It is also interesting to estimate the dynamical time at each position along the propagation streamline based on the derived results for . An illustration of this point is reported in Appendix F.
5 Discussion
As mentioned in the Introduction, different scenarios have been proposed for the interpretation of the recent observations of layered molecular winds. These scenarios are, in principle, also applicable to the HL Tau case. Ellipses in channel maps and transverse PVs, as well as an increase in velocity in longitudinal PVs are predicted by models of magnetized disk winds (Ray & Ferreira, 2021; de Valon et al., 2022), models of wide-angle winds driven swept-up shells (Lee et al., 2000; Zhang et al., 2019), unified models with a jet-bearing wide-angle X-wind (Shang et al., 2023; Ai et al., 2024). On the other hand, flow rotation is expected in both magnetically driven winds, which extract angular momentum from the disk, and thermal photoevaporative winds (Alexander et al., 2014), which conserve their initial angular momentum, but do not slow down the rotation of the disk. The layered wind structure could also be due to the interaction with a pulsed axial jet (e.g., Rabenanahary et al. (2022)). Each of these possibilities deserves a dedicated analysis. In the following, we focus on the scenario of extended MHD disk wind (DW), deferring the comparison of the observed trends with other classes of models to forthcoming papers.
5.1 Basics of the MHD Disk-Wind theory.
According to the Disk-Wind theory, a young star-disk system is embedded in a large-scale hourglass-shaped, twisted magnetic field inherited from the collapse of a magnetized rotating cloud. In this configuration, and provided sufficient ionization, a wind can be launched by the action of magnetic effects from the surface of the rotating disk. The outflow carries away vertically angular momentum, and the consequent braking of the disk rotation allows accretion through the disk to occur. Self-similar steady and axisymmetric DW solutions in ideal MHD conditions have proven to be successful in reproducing the observed properties of the axial atomic jets (see, e.g., Frank et al. 2014; Pascucci et al. 2023). The wind structure predicted by such models, that is, nested rotating gas shells, with higher poloidal velocities along the inner shells, is actually observed also in the outer CO outflows, as shown above for the HL Tau case. However, to properly describe global disk winds from an extended launch region, the models have to include non-ideal effects in the disk. Among them, ambipolar diffusion (AD) between neutrals and ions becomes dominant when ionization is scarce (e.g. Pinto et al. (2008)). Magnetized winds also arise in this case, as described, for instance, in the review by Lesur et al. (2023).
An important quantity of the DW theory is the parameter , which represents the ratio between the extracted and the initial specific angular momentum. Following the definition of Blandford & Payne (1982), can be expressed as:
where is the total specific angular momentum carried away by the MHD wind along a streamline, in the form of both matter rotation and magnetic torsion. is the Keplerian angular rotation speed at the launch point , is the mass volume density and and are the poloidal and toroidal components of the magnetic field in the wind. The higher the value of , the more efficient is the extraction of angular momentum and the acceleration of the wind. The parameter is often referred to as the ’magnetic lever arm’, as it can be expressed as , where is the Alfvén radius, that is, the cylindrical radius at which the poloidal velocity reaches the Alfvén velocity . The locus of points in the wind where this occurs is called Alfvén surface, and it indicates the region above the disk where the increasing inertia of the flow starts to overcome the magnetic push. In self-similar steady DW solutions, this surface has a conical shape.
In the DW scenario is constant along a streamline, which means that the angular momentum is gradually transferred from the magnetic field to the motion of the particles. Continuing along the streamline, the component , representing the normalized specific angular momentum, increases and approaches in the asymptotic regime, in which the magnetic contribution is sensibly reduced. The distance from the disk at which this occurs is larger for higher values of , and can be several orders of magnitude higher than the altitude of the Alfvén surface. However, may never reach , as the magnetic field maintains asymptotically a fraction of the available energy.
The importance of gas pressure versus magnetic effects in the disk can be parametrized by the plasma at the midplane, defined as , where is the sound speed and are the mean density and vertical magnetic field threading the disk. Unfortunately, the magnetic field amplitude in the disk is not available from observations. However, according to theoretical studies for self-similar disk wind solutions, is expected to increase with smaller (e.g. Jacquemin-Ide et al. 2019; Lesur 2021). It follows that the determination of from the observations is crucial not only to identify the properties of the wind, but also to constrain the importance of magnetic effects in the disk.
Self-similar DW solutions from fully turbulent disks have been found for a wide range of values at both high and low (Ferreira, 1997; Casse & Ferreira, 2000a; Jacquemin-Ide et al., 2019; Zimniak et al., 2024). On the other hand, in disk regions where the ionization is quenched in the midplane, as expected in the so-called ’dead zone’ of the disk, solutions are dominated by AD, and low values of are expected, up to about 3 (Lesur, 2021). Recent results show that the dead zone could actually start closer than 1 au from the star in T Tauri disks (Flock et al., 2019), and extend to the outer disk regions of interest here.
In the next Section the principles of the DW theory will be applied to the derivation of relevant quantities in the HL Tau wind, such as the location of origin of the flow in the disk, and the magnetic lever arm. The variation with distance from the source of the wind parameters will also be tested against the prescriptions of these models.
5.2 Derivation of the shell footpoints in the Disk Wind scenario.
Regardless of the conditions adopted for the disk physics, the wind can be treated in ideal MHD. Assuming that the flow is a cold stationary and axisymmetric disk wind, from the measurements described in Sect. 4.1 one can infer the so-called footpoint radius , that is, the radius of the disk annulus from which the wind shell originates. Following the method introduced in Bacciotti et al. (2002) and refined in Anderson et al. (2003), can be found as the real solution of the following third-order equation:
where is the gravitation constant and the mass of the central source. The equation is valid if the enthalpy and the gravitational force exerted by the star are negligible with respect to the kinetic terms at the height above the disk where the wind parameters are measured. These conditions are fulfilled for cold CO molecular flows with poloidal velocities around 10 km sobserved at more than 500 au from a star of solar type, conditions met in our case. In addition, it is assumed that the flow propagates freely. Actually, the environment around HL Tau has a very complex structure, and before reaching the observed location the outflow may have interacted with several elements disturbing the steady propagation, such as accretion streamers (Garufi et al., 2022), winds inflated by nearby stars (Welch et al., 2000), surrounding cavities, and with the lateral wings of the bow-shocks of the inner wind components. However, the regular geometry observed on large scales supports the assumption that the kinematical information survives the occurring perturbations. We note that the observed emission in the CO 2-1 line is seen just because of the density contrast, even if the flow is perfectly laminar, as the emission does not require shock heating to be produced, since the energy of the upper level is only 17 K. However, the presence of external perturbations cannot be excluded a priori, and our simple model should be regarded as a first approximation to the real situation. The investigation of such effects has been deferred to forthcoming work.
The results of the application of Eq. 3 to the features of family B are collected in column (14) of Table 3, and are graphically illustrated in Fig. 11. For each wind shell, the footpoint radii derived from the measurements at different separations from the star turn out to have an almost constant value, with a spread of only a few au.
Interestingly, the footpoint radii of each of the B1a, B1b, and B2 shells cluster around the radius of one of three consecutive rings in the dust emission distribution in the disk, located at 58, 72 and 86 au from the source, respectively (labeled B4, B5 and B6 in ALMA Partnership et al. (2015))555Offsets from the source in ALMA Partnership et al. (2015) have been rescaled to a distance from HL Tau of 147.3 pc, following Galli et al. (2018).. However, note that the uncertainty of 20% in the determined footpoint radius (estimated by testing the sensitivity of the calculation in Eq. 3 to variations in the input parameters) also covers the positions of adjacent disk substructures.
An equivalent derivation of the shell footpoints is illustrated by the diagnostic diagram of Fig. 12, constructed following the procedure described in Ferreira et al. (2006). Under the same assumptions, in fact, it can be shown that and for a given streamline depend only on the two parameters and . The location of the measured () values on the grid of theoretical curves drawn in the diagram gives estimates of the radius of the footpoint r0 and of the parameter . In HL Tau, for each individual shell the points lie along or near a curve at a defined value of , corresponding to the position of one of the disk rings B4, B5, and B6, and progress toward higher values of as the distance from the source increases. As discussed in the next section and in Appendix G, this trend is consistent with the one expected in a DW scenario.
We warn that the positions of the wind footpoints should be compared with the distribution of disk substructures in the gas, rather than in the dust, and these may not coincide. It is well known that bumps in the gas pressure act as dust traps, but the average shift of the peak positions of gas pressure and dust emission depends on the model adopted (see, e.g. Riols et al. 2020; Hu et al. 2022). Yen et al. (2019a) reports profiles of gas emission in the HL Tau disk in the molecular lines of HCO+(3-2) and HCO+(1-0). The profiles are quite shallow, with no clear substructure detected, but it is interesting to note that the footpoint radii of the B-shells are in the vicinity of peaks in the HCO+(3-2)/HCO+(1-0) ratio (see their Fig. 9), which, under the assumption of constant temperature, can be considered proportional to the gas density. According to this comparison, therefore, the wind shells appear to arise from gas rings. However, the association of the footpoint locations with rings remains tentative at this stage, as a more reliable identification would require higher precision in the measurements of the gas tracers.
The derived footpoint radii are larger than our spatial resolution limit ( 40 au). In fact, with reference to Fig. 14, the base of the red-shifted wind appears resolved in the channel maps at low velocities between VLSR= 7.4 and 9.4 km s. However, the emission pattern is not clear in these maps, as contributions at low velocity and close to the star may come both from the outflow base and from parts of the envelope, possibly from accreting streamers, or from some parts of the disk. Nevertheless, as discussed in Appendix A, this velocity interval is consistent with that in which, under the DW hypothesis, we expect to resolve the flow base in these observations. In the channel maps at VLSR km sthe wind base is not resolved on the SE side. At this velocity the flow appears to originate within a region inside the resolution limit of 40 au, it propagates initially with a large opening angle, up to a projected distance of about 0.”5 from the disk, and then it appears to recollimate. This trend is again consistent with the DW scenario, following which the inner shells coming from an unresolved footpoint are subject to a faster acceleration, have an initial large opening angle, which then reduces to the collimated value (cf. Fig. 13, first panel). We could not apply the same analysis to the NW side as the peak of the disk emission at these velocities covers the signal from the wind.
5.3 Variation of parameter in the wind shells
In Section 5.2, we determined at each distance sampled from the source through either its direct calculation as in Footnote 3, or graphically as in the diagram of Fig. 12. For all points relative to HL Tau is within the range = 2.5-4.0, and for each shell, the points relative to PV⟂ in subsequent offsets follow curves at constant , and are found at higher values of as the distance from the source increases. We note that the same behavior is shown by the points corresponding to theoretical DW models, in a zone in which the asymptotic regime has not yet been reached (see Appendix G). This appears to indicate that also in HL Tau we are probing a transition region in which the magnetic field is still transferring angular momentum to the matter. However, the increasing vicinity of the points for higher appears to indicate that the distances probed are not far from the asymptotic regime. This point should be farther away for the outer streamlines than for the inner ones, due to the self-similar nature of the flow (Ferreira et al., 2006). This is in line with the fact that a similar value of is attained closer to the star for the B1 shells than for B2 (sampled at a larger distance).
We also show in Fig. 12 the values of and derived for the shells identified in the HH30 CO outflow (Louvet et al., 2018; López-Vázquez et al., 2024) and in the bright conical DG Tau B CO outflow (de Valon et al., 2022). Interestingly, the and measurements in HH 30 are consistent with a curve at constant , while the points in DG Tau B follow a curve at constant . We caution, however, that the choice of the sampling points of the tomography in DG Tau B differs from ours (see Appendix E). In fact, the behavior determined in DG Tau B is expected in the case in which the tomography is probing a surface of varying in a homogeneous disk wind. Averaging our () values along each HL Tau shell would lead to a trend similar to that observed in DG Tau B, although with larger values of and .
5.4 Comparison with self-similar MHD disk-wind solutions.
The quantity is only one component of the magnetic lever arm. To find observationally the value of the full with the methods described above, one should examine the wind in the shell until the point at which asymptotically reaches . However, this is hampered by the progressive weakness of the signal and by the interaction of the flow with the environment. A possible solution is to compare the variation of the wind parameters in the three shells with the trends prescribed by the DW models calculated for different values of . This also constitutes a test of the validity of the application of the DW theory to the observed outflow.
We do not attempt to explore the full range of possible self-similar solutions, but we show expectations from only three wind models that together encompass the range of values compatible with the observations. Turbulence-dominated disk models should be used for MHD winds anchored below a few au, and AD-dominated models further out. However, in both cases, the MHD wind is described by the same set of ideal MHD equations, and what differs is only the physical conditions at the jet launching point.
We use two solutions, L1 and L2, corresponding to cold disk winds arising from an AD dominated disk (Lesur, 2021), with, respectively, and 63 and corresponding to = 1.56 and = 2.76. Both solutions assume an ambipolar diffusion parameter = 1 and no Ohmic diffusion. The disk and its atmosphere were assumed to be cylindrically isothermal (sound speed ), with a disk aspect ratio = 0.1 (where is the geometric thickness of the disk , with the Keplerian angular velocity). The value achieved by solution L2 is the highest published so far for AD-dominated disks.666The solutions in Lesur (2021) used in our work are freely available at https://github.com/glesur/PPDwind.
In order to investigate the comparison with an MHD wind with a higher magnetic lever arm, we computed a warm solution corresponding to the case of a fully turbulent disk, similar to the one presented in Tabone et al. (2020) to reproduce the ALMA observations of the rotating SO wind in HH212. Our solution N5 describes a wind that arises from a thinner disk with = 0.01, 10, an effective magnetic Prandtl number of 1, a turbulence parameter =2.3, a turbulence anisotropy = 4.2 and no turbulent pressure (=0) (Zimniak et al., 2024). Despite its large magnetic field, the ejection efficiency is high and the resulting outflow is dense. It achieves a magnetic lever arm 5.64 thanks to additional heating at the base of the wind (see the description in Casse & Ferreira 2000b). This heating is localized at the disk surface, and requires only 2% (=0.02, see Eq. 25 in that work) of the dissipated power within the disk. Further up, the wind becomes adiabatic.
The comparison with the models is illustrated in Fig. 13. The points relative to the three B shells, with the same color code as in Fig. 10, describe the wind parameters normalized by their respective average footpoint radius and by the corresponding Keplerian velocity . The bottom curves are from the wind model from the AD-dominated disk region in Lesur (2021), with curve L1 (orange) corresponding to the solution for , =1.56, and curve L2 (blue) for , =2.76. The upper curve N5 (purple) corresponds to the model of wind from a fully turbulent disk region that provides =5.64 (cf. Zimniak et al. 2024).
First, it can be noted that once normalized, the points for the three observed shells in each panel tend to overlap. This is expected for self-similar solutions, in the case in which is similar for all shells. Secondly, the values of the normalized data are in the same range as those predicted by the models. The variation of the quantities with is similar to that of the models in this range, but the gradient in the data points appears to be steeper than in the theoretical curves, for all quantities except .
The top panel shows the comparison with the expected width of the shells. The opening angle of the wind is consistent with that predicted by the model N5 at higher . For the other quantities, the observed points lie in an intermediate position between the curves predicted by the two models L2 and N5, corresponding to an expected asymptotic value of between about 4 and 5 and between 63 and 10.
The fact that the observational points lie between the two types of solution may have interesting implications. For winds coming from disk regions dominated by turbulence, the usual assumptions lead to a limiting anchoring radius of a few au; so in principle our observations, indicating larger footpoint radii, should be better described by the L solutions, corresponding to winds arising from AD-dominated disk regions.
This may indicate that the plasma in the outer regions of the HL Tau disk is smaller (that is, the magnetic effects are more important) than commonly assumed for this zone in protoplanetary disks. A stronger field produces stronger vertical compression and reduced wind density, leading to a larger . However, other effects enter the wind acceleration mechanism, such as the aspect ratio of the disk, the existence of heating at the disk surface, the amount of effective turbulence, the degree of diffusion anisotropy, and the amount of ambipolar diffusion present in the disk. Regarding the last point, the outer region of the disk might possibly present a level of ambipolar diffusion consistently higher than that assumed in solutions of type L, related to an unexpectedly low level of ionization. This might be due to the screening of cosmic rays (the dominant ionization agent at 50 au) by magnetic mirroring from the wind, as proposed in Cleeves et al. (2015) for the TW Hya disk from chemical arguments. Reducing the irradiation of these particles could thus reduce the mass loading on the field lines and increase (Lesur, 2021). An AD-dominated scenario, leading to a laminar-type configuration in the region where rings/gaps are observed, would also be in line with the fact that the rings appear to be remarkably axisymmetric in HL Tau, pointing towards a low level of turbulence in the plane of the disk (Pinte et al., 2016).
In summary, the physical properties of the molecular CO wind from HL Tau appear to be consistent with a description in terms of a magnetized wind originating from a large region of the disk, up to 90 au and possibly beyond. The idea that not only the axial atomic jet originated in the first au from the star, but also the outer wind may be magnetic in nature implies that feedback from flows onto the disk may occur on a large scale, which can have important consequences for models of disk evolution and planet formation. First of all, such winds may be able to sustain accretion along the disk surface at large distances from the star, in regions where the disk is internally weakly ionized and MHD turbulence, and hence viscous accretion, is suppressed. However, other effects may come into play. As discussed in Pascucci et al. (2025), for example, magnetically driven winds can slow the drift of solids in the direction of the star and affect the migration of planetary embryos.
5.5 Implications of the observed shell separation.
Up to this point, the discussion has regarded observational trends along the direction of the wind streamlines, indicating an overall agreement with a DW scenario.
However, the most puzzling evidence that emerges from the data is the inhomogeneity of the wind across its width, as indicated by the presence of distinct nested shells with enhanced emission. This property, not expected in a standard self-similar DW, is actually found in many molecular outflows observed at the high spatial and spectral resolution offered by ALMA (e.g. Zhang et al. 2019; Fernández-López et al. 2020; López-Vázquez et al. 2024; Ai et al. 2024; de Valon et al. 2020, 2022; Omura et al. 2024; Nazari et al. 2024, cf. Introduction).
The analysis of the HL Tau wind offers the advantage of allowing a comparison between the structures in the wind and in the disk. The derived configuration of the wind shells, the large estimated footpoint radii, and the indication of origin of the shells from denser gas rings of the disk appear to support recent models that prescribe the spontaneous formation of a ring-gap substructure in the disk and of a connected inhomogeneous wind, as a result of magnetic instabilities (Béthune et al., 2017; Suriano et al., 2018, 2019; Riols & Lesur, 2019; Riols et al., 2020; Cui & Bai, 2021; Hu et al., 2022). In these numerical studies the secular evolution of protoplanetary disks is studied including ambipolar diffusion in the generalized Ohm’s equation. As a result, one sees the rapid development of a self-organization of the magnetic field in the disk, which accumulates in the regions of lower mass density. In this way a system of rings and gaps forms quickly in the disk, with a greater concentration of magnetic field in the gaps. The final configuration also sees the formation of a transversely inhomogeneous disk wind arranged in nested shells, with alternate layers of enhanced density originating from the gaseous disk rings and layers of faster lighter flow originating from the highly magnetized gaps. The ALMA observations like the ones described in this paper provide maps of the intensity of the line emission, and hence are prone to detect the wind layers of higher density, that is, in the above scenario, the layers arising from the disk rings.
The interest of this approach is the alternative it offers to models in which the disk substructure is due to the presence of young planets. Given the lack of detection of planetary bodies in the HL Tau disk (Mullin et al., 2024) as well as in the other structured disks (with the exception of PDS 70, where, however, the planets reside in a very large cavity, Keppler et al. 2018), the point is of crucial importance, as it indicates that magnetic effects must be taken into account in the investigations aimed at the determination of the timescale of formation of the planets themselves.
However, a number of issues need to be addressed before we can confirm the validity of the proposed association. First of all, the value of the magnetic lever arm assumed in this class of numerical models is generally 2. A similar simulation for , as suggested by our analysis, would be desirable. Secondly, the vertical extension of the structuration of the flow varies among the models, but in general the simulations do not reach the region at 103 au from the star in which we observe the wind structures with ALMA. An extension of numerical studies to large vertical scales would be needed, but it would come at the expense of a consistent computational effort. Finally, in these simulations, the disk structures are often seen to slowly move (Suriano et al., 2019; Martel & Lesur, 2022), questioning the survival of their connection to the wind. However, the motion occurs on secular timescales, suggesting that the coherence of the magnetic flux tubes is probably not disrupted. Clearly, the applicability of the class of models described above to the observed cases needs to be tested with refined theoretical analyses, which, however, are out of the scope of this work.
6 Conclusions
We investigated the properties of the molecular CO wind associated with the young HL Tau system by means of a detailed spectro-imaging analysis of the 12CO (2-1) line emission, observed at 03 angular resolution in the context of the ALMA-DOT campaign (Podio et al., 2020a; Garufi et al., 2021). Our analysis confirms previous results derived from integrated intensity maps, but new insights on the structure of the flow come from the inspection of the individual projections of the Position-Position-Velocity (PPV) ALMA datacube in the brighter and less perturbed red-shifted lobe.
A complex substructure is revealed in the velocity channel maps, constituted by concatenated bubble- and arc-shaped features that progressively increase in size and distance from the source with velocity. The conical shape of the wind appears to be the result of superposition of the limb-brightened borders of these features. The position-velocity (PV) diagram formed along the system axis presents a fan of separated almost linear traces opening from the source region, and with increasing velocity with offset from the star. The traces have strict spatial correspondence with the arcs in the channel maps. Finally, PV diagrams taken across the flow at stepped separations from the source show a series of concatenated ellipse-shaped features that correspond directly to the traces in the longitudinal PV diagram and in the channel maps. The ellipses are tilted with respect to the system axis, indicating rotation of the flow in clockwise direction looking from the SW lobe toward the star, which matches the disk rotation sense.
On the basis of these projections we infer a global structure of the emission in the PPV datacube consisting of a series of distinct concatenated paraboloid-like surfaces with the apex toward the source, with progressively larger aperture as the angle of their axis with respect to the velocity axis increases (cf. Fig. 7). In the 3D physical space this translates into a finite number of coaxial, distinct, rotating nested gas shells. Ellipse fitting of the three directly identifiable structures in adjacent transverse PV diagrams allows a tomographic reconstruction of the morphology and kinematics of the outer three shells of the flow. We find smaller radii, higher poloidal and toroidal velocities, and steeper increases of the velocity with distance from the source for shells progressively closer to the system axis.
This structure can be reproduced by a number of different wind acceleration mechanisms. In this paper, we focus on an analysis of the applicability of magnetohydrodynamic (MHD) disk winds (DW) accelerated from an extended region of the disk surface. We defer the comparison with other classes of models to future investigations.
Under the DW hypothesis, we derive from the rotation properties of the flow that each examined shell is rooted in the disk within a few au of one of three adjacent disk rings, seen in the continuum emission of the dust, located at 58, 72 and 86 au from the star (ALMA Partnership et al., 2015). However, we caution that, because of the faintness of the examined flow substructures, the uncertainty on the footpoint radii is greater than the distance of the adjacent rings and gaps.
We attempted a determination of the magnetic lever arm in the wind by comparing the observed trends of the wind parameters along the shells with the predictions of models of magnetized winds from either a laminar AD-dominated disk or a fully turbulent disk. The characteristics of the HL Tau wind turn out to be intermediate between the two types of solutions, with a value of . We discuss this finding in the context of the characteristics of the examined models. The retrieved launching region between 50 and 90 au would point to a wind from an AD-dominated outer disk, but solutions with such a high value of do not exist and have yet to be found (current solutions reach = 2.76). A possibility would be to consider solutions for a higher level of AD than is usually assumed, which may be applicable if the wind itself screens the incoming ionizing cosmic rays by magnetic mirroring (Cleeves et al., 2015). An increased field amplitude in the outer disk region with respect to the one commonly adopted could also help justify the high value of , as the ejection efficiency in the wind would be higher.
A magnetically driven disk wind with the above characteristics could sustain disk accretion at large scale by removing angular momentum in regions where the effective turbulence is suppressed because of weak internal ionization. In addition, the presence of distinct denser nested wind shells rooted in disk rings appears to support the results of recent non-ideal MHD numerical simulations for AD-dominated disks, which see the spontaneous formation of a ring-gap substructure in the disk connected with an inhomogeneous layered flow (e.g. Suriano et al. 2019; Riols et al. 2020), in a way alternative to the action of the yet elusive protoplanets. These aspects may have profound implications for the evolution of the disk and the process of planet formation.
The inferred properties of the CO wind call for further analyses of both observational and theoretical type. An analogous scrutiny of the fainter components and substructures in both the red- and blue-shifted lobes in the same datacube, as well as the challenging estimate of the mass and angular momentum transported in each shell, requires the development of dedicated procedures, which is currently underway (Nony et al., in prep.). On the theoretical side, it would be interesting to consider the extension of the numerical simulations of inhomogeneous layered disk winds to the large scales probed by our observations. Then, a comparison of the retrieved results with other classes of wind models is to be attempted, and it will the subject of future works. Finally, new observational studies for this and other systems, currently in progress, will allow us to analyze the dynamic relationship between the outer CO outflow, the coaxial H2 wind and the atomic jet.
These investigations are expected to provide further insight on the nature of the ejection phenomenon and finally to make us able to quantify the feedback of the winds on the disk evolution.
Acknowledgements.
This paper uses ALMA data 2018.1.01037.S (PI L. Podio). ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. We thank the anonymous referee for her/his precious comments that led to improved results. The authors are grateful to G. Lesur for fruitful discussions and for his support in using the self-similar MHD disk wind solutions that he has made available to the public at: https://github.com/glesur/PPDwind. The authors thank Z.-Y. Li, Y. Bai, T. Downes, F. Louvet, D. Fedele, and L. Testi for their helpful comments, and K. Rygl, V. Delabrosse, and U. Locatelli for their help with data processing and analysis. This work was supported by Large Grant INAF 2022 ’YSOs Outflows, Disks, and Accretion: towards a global framework for the evolution of planet-forming systems (YODA)’. Part of the research activities were carried out with the contribution of the Next Generation EU funds within the National Recovery and Resilience Plan (PNRR), Mission 4 - Education and Research, Component 2 - From Research to Business (M4C2), Investment Line 3.1 - Strengthening and creation of Research Infrastructures, Project IR0000034 – ”STILES - Strengthening the Italian Leadership in ELT and SKA”. LP and CC acknowledge the projects PRIN-MUR 2020 - ”BEYOND-2p: Astrochemistry beyond the second period elements” (Prot. 2020AFB3FX); ASI-Astrobiologia 2023 - ”MIGLIORA-Modeling Chemical Complexity”, (F83C23000800005); INAF-GO 2023 - ”PROTO-SKA-Exploiting ALMA data to study planet-forming disks: preparing the advent of SKA” (C13C23000770005); INAF MiniGrant 2022 - ”Chemical Origins” (PI: L. Podio); National Recovery and Resilience Plan (NRRP), Mission 4, Component 2, Investment 1.1, Call for tender No. 104 - Next Generation EU – 2022JC2Y93 ”Chemical Origins: linking the fossil composition of the Solar System with the chemistry of protoplanetary disks” (CUP J53D23001600006) - Grant Assignment Decree No. 962 adopted on 30.06.2023 by the Italian Ministry of University and Research.References
- Ai et al. (2024) Ai, T.-H., Liu, C.-F., Shang, H., Johnstone, D., & Krasnopolsky, R. 2024, ApJ, 964, 147
- Alexander et al. (2014) Alexander, R., Pascucci, I., Andrews, S., Armitage, P., & Cieza, L. 2014, in Protostars and Planets VI, ed. H. Beuther, R. S. Klessen, C. P. Dullemond, & T. Henning, 475–496
- ALMA Partnership et al. (2015) ALMA Partnership, Brogan, C. L., Pérez, L. M., et al. 2015, ApJ, 808, L3
- Anderson et al. (2003) Anderson, J. M., Li, Z.-Y., Krasnopolsky, R., & Blandford, R. D. 2003, ApJ, 590, L107
- Andrews et al. (2018) Andrews, S. M., Huang, J., Pérez, L. M., et al. 2018, ApJ, 869, L41
- Anglada et al. (2007) Anglada, G., López, R., Estalella, R., et al. 2007, AJ, 133, 2799
- Bacciotti et al. (2002) Bacciotti, F., Ray, T. P., Mundt, R., Eislöffel, J., & Solf, J. 2002, ApJ, 576, 222
- Beck et al. (2008) Beck, T. L., McGregor, P. J., Takami, M., & Pyo, T.-S. 2008, ApJ, 676, 472
- Béthune et al. (2017) Béthune, W., Lesur, G., & Ferreira, J. 2017, A&A, 600, A75
- Bjerkeli et al. (2016) Bjerkeli, P., van der Wiel, M. H. D., Harsono, D., Ramsey, J. P., & Jørgensen, J. K. 2016, Nature, 540, 406
- Blandford & Payne (1982) Blandford, R. D. & Payne, D. G. 1982, MNRAS, 199, 883
- Casse & Ferreira (2000a) Casse, F. & Ferreira, J. 2000a, A&A, 353, 1115
- Casse & Ferreira (2000b) Casse, F. & Ferreira, J. 2000b, A&A, 361, 1178
- Cleeves et al. (2015) Cleeves, L. I., Bergin, E. A., Qi, C., Adams, F. C., & Öberg, K. I. 2015, ApJ, 799, 204
- Codella et al. (2020) Codella, C., Podio, L., Garufi, A., et al. 2020, A&A, 644, A120
- Cui & Bai (2021) Cui, C. & Bai, X.-N. 2021, MNRAS, 507, 1106
- De Valon (2021) De Valon, A. 2021, PhD thesis, Universite de Grenoble Alpes
- de Valon et al. (2020) de Valon, A., Dougados, C., Cabrit, S., et al. 2020, A&A, 634, L12
- de Valon et al. (2022) de Valon, A., Dougados, C., Cabrit, S., et al. 2022, A&A, 668, A78
- Delabrosse et al. (2024) Delabrosse, V., Dougados, C., Cabrit, S., et al. 2024, A&A, 688, A173
- Fernández-López et al. (2020) Fernández-López, M., Zapata, L. A., Rodríguez, L. F., et al. 2020, AJ, 159, 171
- Ferreira (1997) Ferreira, J. 1997, A&A, 319, 340
- Ferreira et al. (2006) Ferreira, J., Dougados, C., & Cabrit, S. 2006, A&A, 453, 785
- Flock et al. (2019) Flock, M., Turner, N. J., Mulders, G. D., et al. 2019, A&A, 630, A147
- Frank et al. (2014) Frank, A., Ray, T. P., Cabrit, S., et al. 2014, in Protostars and Planets VI, ed. H. Beuther, R. S. Klessen, C. P. Dullemond, & T. Henning, 451–474
- Galli et al. (2018) Galli, P. A. B., Loinard, L., Ortiz-Léon, G. N., et al. 2018, ApJ, 859, 33
- Garufi et al. (2021) Garufi, A., Podio, L., Codella, C., et al. 2021, A&A, 645, A145
- Garufi et al. (2020) Garufi, A., Podio, L., Codella, C., et al. 2020, A&A, 636, A65
- Garufi et al. (2022) Garufi, A., Podio, L., Codella, C., et al. 2022, A&A, 658, A104
- Hu et al. (2022) Hu, X., Li, Z.-Y., Zhu, Z., & Yang, C.-C. 2022, MNRAS, 516, 2006
- Jacquemin-Ide et al. (2019) Jacquemin-Ide, J., Ferreira, J., & Lesur, G. 2019, MNRAS, 490, 3112
- Keppler et al. (2018) Keppler, M., Benisty, M., Müller, A., et al. 2018, A&A, 617, A44
- Klaassen et al. (2016) Klaassen, P. D., Mottram, J. C., Maud, L. T., & Juhasz, A. 2016, MNRAS, 460, 627
- Lee et al. (2017) Lee, C.-F., Ho, P. T. P., Li, Z.-Y., et al. 2017, Nature Astronomy, 1, 0152
- Lee et al. (2018) Lee, C.-F., Li, Z.-Y., Codella, C., et al. 2018, ApJ, 856, 14
- Lee et al. (2000) Lee, C.-F., Mundy, L. G., Reipurth, B., Ostriker, E. C., & Stone, J. M. 2000, ApJ, 542, 925
- Lesur et al. (2023) Lesur, G., Flock, M., Ercolano, B., et al. 2023, in Astronomical Society of the Pacific Conference Series, Vol. 534, Protostars and Planets VII, ed. S. Inutsuka, Y. Aikawa, T. Muto, K. Tomida, & M. Tamura, 465
- Lesur (2021) Lesur, G. R. J. 2021, A&A, 650, A35
- Lodato et al. (2019) Lodato, G., Dipierro, G., Ragusa, E., et al. 2019, MNRAS, 486, 453
- López-Vázquez et al. (2024) López-Vázquez, J. A., Lee, C.-F., Fernández-López, M., et al. 2024, ApJ, 962, 28
- Louvet et al. (2018) Louvet, F., Dougados, C., Cabrit, S., et al. 2018, A&A, 618, A120
- Lumbreras & Zapata (2014) Lumbreras, A. M. & Zapata, L. A. 2014, AJ, 147, 72
- Martel & Lesur (2022) Martel, É. & Lesur, G. 2022, A&A, 667, A17
- McMullin et al. (2007) McMullin, J. P., Waters, B., Schiebel, D., Young, W., & Golap, K. 2007, in Astronomical Society of the Pacific Conference Series, Vol. 376, Astronomical Data Analysis Software and Systems XVI, ed. R. A. Shaw, F. Hill, & D. J. Bell, 127
- Movsessian et al. (2012) Movsessian, T. A., Magakian, T. Y., & Moiseev, A. V. 2012, A&A, 541, A16
- Müller et al. (2001) Müller, H. S. P., Thorwirth, S., Roth, D. A., & Winnewisser, G. 2001, A&A, 370, L49
- Mullin et al. (2024) Mullin, C., Dong, R., Leisenring, J., et al. 2024, AJ, 167, 183
- Mundt et al. (1990) Mundt, R., Buehrke, T., Solf, J., Ray, T. P., & Raga, A. C. 1990, A&A, 232, 37
- Nazari et al. (2024) Nazari, P., Tabone, B., Ahmadi, A., et al. 2024, A&A, 686, A201
- Nisini et al. (2024) Nisini, B., Navarro, M. G., Giannini, T., et al. 2024, ApJ, 967, 168
- Omura et al. (2024) Omura, M., Tokuda, K., & Machida, M. N. 2024, ApJ, 963, 72
- Pascucci et al. (2025) Pascucci, I., Beck, T. L., Cabrit, S., et al. 2025, Nature Astronomy, 9, 81
- Pascucci et al. (2023) Pascucci, I., Cabrit, S., Edwards, S., et al. 2023, in Astronomical Society of the Pacific Conference Series, Vol. 534, Protostars and Planets VII, ed. S. Inutsuka, Y. Aikawa, T. Muto, K. Tomida, & M. Tamura, 567
- Pinte et al. (2016) Pinte, C., Dent, W. R. F., Ménard, F., et al. 2016, ApJ, 816, 25
- Pinto et al. (2008) Pinto, C., Galli, D., & Bacciotti, F. 2008, A&A, 484, 1
- Podio et al. (2019) Podio, L., Bacciotti, F., Fedele, D., et al. 2019, A&A, 623, L6
- Podio et al. (2011) Podio, L., Eislöffel, J., Melnikov, S., Hodapp, K. W., & Bacciotti, F. 2011, A&A, 527, A13
- Podio et al. (2020a) Podio, L., Garufi, A., Codella, C., et al. 2020a, A&A, 642, L7
- Podio et al. (2020b) Podio, L., Garufi, A., Codella, C., et al. 2020b, A&A, 644, A119
- Podio et al. (2021) Podio, L., Tabone, B., Codella, C., et al. 2021, A&A, 648, A45
- Pyo et al. (2006) Pyo, T.-S., Hayashi, M., Kobayashi, N., et al. 2006, ApJ, 649, 836
- Rabenanahary et al. (2022) Rabenanahary, M., Cabrit, S., Meliani, Z., & Pineau des Forêts, G. 2022, A&A, 664, A118
- Ray & Ferreira (2021) Ray, T. P. & Ferreira, J. 2021, New A Rev., 93, 101615
- Riols & Lesur (2019) Riols, A. & Lesur, G. 2019, A&A, 625, A108
- Riols et al. (2020) Riols, A., Lesur, G., & Menard, F. 2020, A&A, 639, A95
- Segura-Cox et al. (2020) Segura-Cox, D. M., Schmiedeke, A., Pineda, J. E., et al. 2020, Nature, 586, 228
- Shang et al. (2020) Shang, H., Krasnopolsky, R., Liu, C.-F., & Wang, L.-Y. 2020, ApJ, 905, 116
- Shang et al. (2023) Shang, H., Liu, C.-F., Krasnopolsky, R., & Wang, L.-Y. 2023, ApJ, 944, 230
- Sheehan & Eisner (2017) Sheehan, P. D. & Eisner, J. A. 2017, ApJ, 840, L12
- Sheehan & Eisner (2018) Sheehan, P. D. & Eisner, J. A. 2018, ApJ, 857, 18
- Stephens et al. (2023) Stephens, I. W., Lin, Z.-Y. D., Fernández-López, M., et al. 2023, Nature, 623, 705
- Suriano et al. (2018) Suriano, S. S., Li, Z.-Y., Krasnopolsky, R., & Shang, H. 2018, MNRAS, 477, 1239
- Suriano et al. (2019) Suriano, S. S., Li, Z.-Y., Krasnopolsky, R., Suzuki, T. K., & Shang, H. 2019, MNRAS, 484, 107
- Tabone et al. (2017) Tabone, B., Cabrit, S., Bianchi, E., et al. 2017, A&A, 607, L6
- Tabone et al. (2020) Tabone, B., Cabrit, S., Pineau des Forêts, G., et al. 2020, A&A, 640, A82
- Tabone et al. (2018) Tabone, B., Raga, A., Cabrit, S., & Pineau des Forêts, G. 2018, A&A, 614, A119
- Takami et al. (2007) Takami, M., Beck, T. L., Pyo, T.-S., McGregor, P., & Davis, C. 2007, ApJ, 670, L33
- Welch et al. (2000) Welch, W. J., Hartmann, L., Helfer, T., & Briceño, C. 2000, ApJ, 540, 362
- Yen et al. (2019a) Yen, H.-W., Gu, P.-G., Hirano, N., et al. 2019a, ApJ, 880, 69
- Yen et al. (2019b) Yen, H.-W., Takakuwa, S., Gu, P.-G., et al. 2019b, A&A, 623, A96
- Zapata et al. (2015) Zapata, L. A., Lizano, S., Rodríguez, L. F., et al. 2015, ApJ, 798, 131
- Zhang et al. (2016) Zhang, Y., Arce, H. G., Mardones, D., et al. 2016, ApJ, 832, 158
- Zhang et al. (2019) Zhang, Y., Arce, H. G., Mardones, D., et al. 2019, ApJ, 883, 1
- Zimniak et al. (2024) Zimniak, N., Ferreira, J., & Jacquemin-Ide, J. 2024, A&A, 692, A99
Appendix A Atlas of velocity channel maps
Figures 14 and 15 present an atlas of channel maps of CO (2-1) emission at VLSR from -3.0 km sto +28.6 km s, in steps of 0.4 km s. The panels at VLSR from +3.0 to +8.6 km sare centered on the source, to visualize the low-velocity emission in both lobes, while outside this range, emission is detected only in either the blue- or in the red-shifted lobe, so the maps cover the relevant quadrant, with the source positioned in one corner.
The emission in the blue-shifted NE lobe is very faint. It consists of a large arc-shaped feature opening toward the NE seen at progressively larger distances from the source with increasing blue-shifted velocities. A finger-like feature is identified at almost all velocities in about the same position north to the source and appears to be independent on the moving arc.
The red-shifted SW lobe is distributed in two main components, a low-velocity wide-angle wind and a faster inner flow of apparent conical shape, with an opening angle decreasing with velocity. To quantify the variation of the aperture, we determined, for selected channels from 9 to 28 km sand separately for the SE and NW sides, the angle between the system axis and a line connecting the source and the position of the peak emission at the flow borders (identified by the contour profiles), at an offset of 45 and 75 from the source. The results of this analysis are shown in Figure 16. The half-opening angle decreases on average with velocity, from about 40 2 at 10 km sto 27 2 at 26 km s. The difference in value displayed in a given channel by the various curves reflects the deviations in the shape of the flow from a perfect cone.
Inscribed in the apparent cones one finds a series of arcs and bubble-like features. Each of these features can be followed over a number of adjacent channels, and it maintains the same shape and orientation with respect to the axis. The distance from the source and the size of each feature increase monotonically with VLSR. As an example, Fig. 17 illustrates the shift in position of the arc-shaped structures of family B with increasing velocity.
The base of the red-shifted wind (with reference to the SE side with respect to the axis, as the NW side hosts the emission from the receding part of the disk) appears resolved in the channel maps at low velocities between VLSR= 7.4 and 9.4 km s. Although the emission pattern is not clear, this velocity interval is consistent with the one expected under the DW hypothesis. In this scenario, the flow starts with at the footpoint (corresponding to VLSR= Vsys), and it is gradually faster as it leaves the disk. However, due to the absorption of the ambient medium, no material with VLSR 7.3 km s is visible (see, e.g. Fig. 5). On the other hand, the maximum VLSR at which the wind base will be imaged is VLSR= Vsys+ , where is the poloidal velocity achieved at an altitude 60 au above the disk. Compared with the curves of the DW model discussed in Section 5.4, is expected to be between 2.0 and 3.5 km sfrom B2 to B1a (see Fig. 13, where the observed quantities are normalized by the corresponding and ). Projecting along the l.o.s. and adding Vsys, one finds that the base of the flow should be visible in channel maps up to VLSR 8.4 – 9.5 km s, as we find, with the highest VLSR corresponding to inner B1a. As predicted by the model, in the channel maps at VLSR km sthe wind base is not resolved.
Appendix B Atlas of transverse Position-Velocity diagrams and application of the ellipse fitting
Figures 18, 19 and 20 present an atlas of the transverse Position-Velocity diagrams (PV⟂) obtained for the red lobe with a 03-wide virtual slit centered on the axis and oriented at PA=318∘ (i.e. perpendicular to the axis and with positive offsets toward NW), at separations from the source from 35 to 10″ in steps of 05
The left panels show the observed emission in grayscale, while the right panels illustrate the same emission with a contour map. Superposed to the latter, for the PV cuts from 45 to 10″, mathematical ellipse fits of the features are reported. The fit is performed with the purpose of identifying the lateral edges of the observed features. The coordinates of these points are used in a tomographic reconstruction of the morphology and kinematics of the flow shell (see Section 4.1).
The critical aspect of the procedure is the difficulty in identifying unambiguously the traces corresponding to each shells B1a, B1b, and B2 in the faint or crowded regions of the transverse PV diagrams. To give an identification as reliable as possible, a cross-check on the positions of the emission peaks common to different projections of the datacube was performed. Such a correspondence is expected if the observed features originate from cuts of coherent structures in the datacube. To this aim, we first found sequences of relative intensity peaks in each diagram that could trace a feature, all above an emission threshold of a 3 level, where mJy beam-1 is the average over all panels estimated in a region along the axis at VLSR 20 km s, where the signal of the traces is almost absent. We then looked for the signature of these sequences in the other two projections of the data cube. In particular, the position of the traces in the fan of the PV∥ diagram (Fig. 5) identifies the expected velocity of each feature along the axis, while in the channel maps, the position of the peaks at the intersection of the arc-like features with the virtual perpendicular cut at a distance from the source provides information on the expected offset of the emission in the PV⟂ diagram of offset at the velocity of the channel map. The quantitative comparison of the coordinates in the three projections allowed the selection of the sequence of peaks belonging to a single feature in each diagram.
Then, a mathematical fitting routine based on a least-squares algorithm was applied to each identified sequence of relative intensity peaks. The fitting is anchored to points in the upper arc and side edges, but not in the lower arc, as the latter is difficult to identify due to overlap of low-velocity traces, signal absorption, and deformation due to flow inclination. The appearance of the traces in the PV⟂ projections suggests an ellipse as the geometrical form closest to the actual shape of the features, so the fit requires a number of input points greater than five. Sampling sets that met the above criteria could be found for shells B1a and B1b at offsets from 45 to 7″from the star, and for B2 from 6″to 10″. The input points are reported as crosses in Figures 18 - 20, with the fitted ellipses drawn in the corresponding color. The routine provides the coordinates of the center, the length of the semi-major and semi-minor axes, and the inclination with respect to the horizontal direction (i.e. VLSR= constant).
The SE and NW edges of the fitted ellipses, that is, the points on the curve farthest from the axis, are considered the vertices of each feature at the borders of the shell (in a couple of cases where the geometric extreme of the ellipse had no emission associated, the closest emission peak was selected). The coordinates of these points provide the quantities necessary for the tomographic reconstruction.
With respect to the error estimate, statistically, the applied fitting procedure is expected to reduce the uncertainty with respect to that of the input points. In practice, however, the advantage is contrasted by the difficulty in identifying the traces, the lack of reliable sampling points in the lower part of the curve, the difference of the actual trace shape with respect to a geometrical ellipse, and the absence of emission at the extremes in a few cases. Therefore, we consider the uncertainty in the final coordinate values comparable to that of the input points in the spectral images, that is, in position and km s in velocity, reported as a small magenta ellipse in the right panels.
Appendix C Influence of wiggling and precession
We tested the influence of wiggling and precession on the derivations in Section 4.1 by analyzing the deviation of the centers of the fitted ellipses of features B with respect to the axis of the PV⟂ diagrams (the perpendicular to the virtual slit of the diagram), corresponding in the sky to PA = 228°. The axis of propagation of the shells is found to deviate slightly from this direction, by 0.5, 2 and 3 degrees toward SE, for the B1a, B1b, and B2 shells, respectively. The scatter of the centers with respect to these axes shows a hint of wiggling pattern consistent among the shells, but the shift is anywhere lower than 3/5 of the observational uncertainty on the shell radius. In any case, our derivation of , Vz and Vϕ relies on the relative positions of the apexes of the ellipses at the borders of the flow and does not use the information on the position of the ellipse center. Therefore, the determination of these quantities is not affected by wiggling, as long as the shell retains its coherence. In contrast, the different PA of the flow with respect to the perpendicular to the virtual slit can introduce a bias. We checked the importance of this effect by evaluating the expected variation in the worst case, that is, for shell B2 at a large offset from the star. The difference between the measured and real values of and is negligible, while the measured turns out to be larger by 1% of the real one, and smaller by 8%. In any case, these values are lower than the observational uncertainty 5% and 10% in the poloidal and toroidal velocity, respectively, which justifies the adoption of Eq. 1 for our analysis.
Appendix D Estimate of the radial velocity V
In Section 4.1 the assumption is made that the wind flows along the detected shells. Formally, this translates into the fact that in the cylindrical coordinate system aligned with the red-shifted flow axis and centered on the star, the poloidal velocity is tangent to the shell surface. To test the validity of this assumption, one should verify that ), where is the local tangent angle to the shell, determined from the derivative of the measured and reported in Footnote 3. While and at a given offset from the source are found directly from the measurements at the lateral apexes of the fitted ellipses in one single diagram, the search for is complicated by the combination of the detected variation of along the shell and the inclination of the flow axis. With reference to Fig. 21, and assuming axisymmetry, to determine the radial velocity corresponding to the axial velocity evaluated at offset , one has to measure VLSR at either point A or point B of the same shell section (the blue circle in the figure), that is, at the points where the projection of along the l.o.s. vanishes:
However, points A and B fall at offsets and . Unfortunately, with the given geometry, these points correspond in almost all cases to regions of the diagrams with faint or confused emission, out of the field of view, hidden by the brightness of the gas close to the source or lost because of the CO auto-absorption around Vsys. These factors make the determination of extremely difficult.
In practice, this quantity can be reasonably determined only for shell B2 in the two positions at 95 and 10″, using the cleaner diagram PV∥ and limiting the search to the upper point of type B. Taking into account the radius of the shell in these positions, the searched point B is at offsets 49 and 53, where one finds VLSR(B)= 15.5 0.3 km sand 15.7 0.3 km s, respectively. Equation 4 then leads to Vr = 2.6 0.4 km sand 2.8 0.4 km s. In turn, the orientation angle of Vp with respect to the axis turns out to be 15.2° 3.0° at offset 95 and 16.3° 3.2° at offset 10″. According to Footnote 3, the tangent angle at these positions is 17.5°, which implies that here Vp is compatible with being tangential to the shell surface within the errors.
Appendix E Comparison with the tomographic procedure in de Valon et al. (2022).
A tomographic analysis similar to the one presented in Section 4.1 was applied in de Valon et al. (2022) to derive the morphology and the kinematic properties of the molecular outflow from the young star DG Tau B. In this source the flow presents a complex system of internal substructures with characteristics analogous to those of the flow from HL Tau. However, the adopted technique is not identical and leads to different outcomes.
The system of Eq. 1 is equivalent to the one in Eqs. (4)-(7) in Sect. 3 of de Valon et al. (2022)777the inclination angle of the flow with respect to the line of sight differs by 180°with respect to ours, which introduces a minus sign in . In both cases, and are derived from the values of VLSR extracted along the bell profile in the PV⟂ diagrams. The difference resides in the selection of the relevant points, a choice that derives from a different assumption regarding the distribution of the gas in the flow. In de Valon et al. (2022) the wind is considered homogeneous in the transverse direction to the axis. Thus, the points are selected with continuity throughout the bell profile. Under the assumption of a perfectly symmetric flow with respect to the axis, the procedure gives at each distance from the source two continuous functions =VLSR and =VLSR for varying offset . From these, maps of and in the plane are derived over a transversely homogeneous wind (see Fig. 3 in de Valon et al. (2022)). The presence of the observed arcs and cusps in the channel maps and of the ellipse-shaped traces in the PV diagrams does not enter the analysis at this stage. In fact, the features are interpreted as emission enhancements moving in a homogeneous flow. The justification of such enhancements invokes the presence of perturbations in the flow caused by a non-stationary mechanism, influencing either the accretion flow through the disk or the ejection of the axial jet (see Fig. 19 in de Valon et al. (2022)).
In our work, the substructures are interpreted as the signature of a wind that is transversely inhomogeneous and structured in a finite number of distinct nested shells. In this way, the observed features naturally arise from the cuts of the wind distribution in the projections of the PPV datacube (cf. Fig. 7). As a consequence, non-stationary mechanisms are not needed to justify the observations. Technically, in our analysis we search for the edges of each ellipse-shaped trace in subsequent PV⟂ diagrams, obtaining for each shell a set of pairs of correlated points (,VLSR-; ,VLSR+) as a function of the distance from the source. From this set, we derive the properties of the flow along that particular shell, and the bell profile is interpreted as the superposition of the limb-brightened traces of the subsequent shells. Part of these traces are bright enough to be clearly identified also inside the bell profile (as the B family), part of them are too faint for a direct analysis and will be examined in a forthcoming work.
Appendix F Evaluation of the dynamical time of the substructures
In the top panel of Fig. 22 we report the dynamical time estimated as , as a function of the distance from the source , for the three substructures analyzed. The dynamical time is higher for the slower outer shells, and for each shell increases monotonically with distance from the source, from 4.8 to 6 102 years for B1a and from 6.0 to 7.1 102 years for B1b, over a distance from 0.9 to 1.4 103 au from the source. For the outer shell B2 is measured to vary between 7.6 and 10 102 years over a distance of 1.2 to 1.9 103 au from the source. The observed trends are compatible with a steady ejection of material and steady propagation along the nested shells as postulated in a magnetohydrodynamic disk wind. The applicability of the disk wind scenario is quantitatively confirmed by Fig. 22, bottom panel, in which the dynamical times derived for the three shells, normalized through their footpoint radii and corresponding Keplerian velocities, are compared to the analogous quantity estimated for the DW models L2 and N5 (cf. Section 5.4). The points show the same trend as the curves, closely following the one for model N5.
Alternatively, assuming that the observed features are swept-up shells caused by subsequent pulsed ejections, from the comparison of at the same altitude one could estimate the time separation between subsequent ejections, obtaining intervals of about one hundred years. However, in this scenario one would rather expect for each shell a constant dynamical time along , as the bulk of the material ejected in each event would reach the current altitude at the same time. However, it cannot be excluded that the interaction with the environment and/or with the material ejected previously causes a slowdown of the flow in the forward part of the shell, leading to an apparent increase in dynamical time at high .
Appendix G Comparison with models - variation of along the streamline
Figure 23 illustrates the similarity of the trend of variation of the specific angular momentum along a streamline. Models L2 (Lesur 2021) and N5 (Zimniak et al. 2024) are calculated for =1 au and a star of one solar mass. The points for shell B2 are normalized by the radius of the retrieved footpoint and by the corresponding Keplerian velocity for a stellar mass of 2.1 M⊙. In this way the wind streamline originates in all cases at 1 au, where it is usual to assume that the turbulence starts, so as to have the most effective comparison between the turbulence- and AD-dominated disk wind models (N and L types of solutions, respectively). In all cases the values are seen to follow closely the curve at constant =1, with a progression toward higher with distance, as expected in steady MHD disk winds.